Please help me solve x^4 - 9 ?
Found 2 solutions by Grinnell, richwmiller:
Answer by Grinnell(63)

How can we put this in familiar form?
9 and 4 are perfect squares. These are hints at something.
we know a^2-b^2 = (a+b)(a-b)
we can do this! x^2^2-3^2=(x^2+3)(x^2-3)=0
3^2+3=0, x=Square root of 3i
3^2-3=0, x=Square root of 3
This is not as complicated as it looks, but you must use the common products of Algebra AND be on the look-out for perfect squares either in the exponents or the real numbers themselves.
square root of 3i AND the square root of 3 are the answers.

Answer by richwmiller(17219)

To be solved there must be an equal sign.
Otherwise it can be factored or simplified but not solved.
let u^2=x^4 so u=x^2
u^2-9=
(u-3)*(u+3)
(x^2-3)*(x^2+3)